OpenStudy (anonymous):

Express the quantity sin^6(x) in a form that does not involve powers of trigonometric functions.

OpenStudy (asnaseer):

Make use of the double angle formula for cos:\[\cos(2x)=1-2\sin^2(x)\]If you rearrange this you will get:\[\sin^2(x)=\frac{1-\cos(2x)}{2}\]

OpenStudy (asnaseer):

Use this and you will end up with the desired answer

OpenStudy (mathmate):

additional hint: similarly \(\large \cos^2(x)=\frac{1+\cos(2x)}{2}\)